# OccamVTS: Distilling Vision Models to 1% Parameters for Time Series Forecasting

### v1

@@ Abstract
Time series forecasting is fundamental to diverse applications, with recent approaches leverage large vision models (LVMs) to capture temporal patterns through visual representations. We reveal that while vision models enhance forecasting performance, 99% of their parameters are unnecessary for time series tasks. Through cross-modal analysis, we find that time series align with low-level textural features but not high-level semantics, which can impair forecasting accuracy. We propose OccamVTS, a knowledge distillation framework that extracts only the essential 1% of predictive information from LVMs into lightweight networks. Using pre-trained LVMs as privileged teachers, OccamVTS employs pyramid-style feature alignment combined with correlation and feature distillation to transfer beneficial patterns while filtering out semantic noise. Counterintuitively, this aggressive parameter reduction improves accuracy by eliminating overfitting to irrelevant visual features while preserving essential temporal patterns. Extensive experiments across multiple benchmark datasets demonstrate that OccamVTS consistently achieves state-of-the-art performance with only 1% of the original parameters, particularly excelling in few-shot and zero-shot scenarios.
@@ 1 Introduction
Time series forecasting (TSF) is a fundamental task in machine learning, underpinning a wide range of critical applications, including energy demand prediction, financial market analysis, meteorological modeling, and traffic flow optimization (Idrees et al. 2019; Kiyasseh et al. 2021; Xu et al. 2021; Bi et al. 2023). The objective of TSF is to anticipate future values based on historical observations of one or more temporally evolving variables. Despite its broad utility, TSF presents significant challenges due to inherent characteristics such as non-stationarity, long-range dependencies, stochastic noise, and the simultaneous presence of localized patterns and global trends across multiple temporal scales.
@@
Deep learning has emerged as the dominant paradigm for TSF, with state-of-the-art models consistently achieving strong performance across diverse benchmarks. These methods operate directly on numerical sequences or their frequency-domain counterparts, enabling rich representations of complex temporal dynamics. Convolutional neural networks (CNNs) and Transformers, for example, are particularly adept at capturing both short-term fluctuations and long-term dependencies (Ismail Fawaz et al. 2019); advanced models such as Autoformer (Wu et al. 2021) and FEDformer (Zhou et al. 2022) explicitly decompose sequences into trend and seasonal components (Bai et al. 2018; Hatami et al. 2018; Li et al. 2019; Liu et al. 2022a; Wen et al. 2023). Distinct from other modalities that rely on abstract semantics, these architectures effectively model time series as structured numerical signals, ultimately grounding predictions in precise and interpretable temporal behaviors.
@@
In parallel, vision models have recently gained significant traction as an alternative and intriguing approach to TSF, driven by the observation that humans can often see meaningful patterns in time series plots. For instance, a seasoned bond trader may visually recognize a shift in trend or momentum in a price curve and form predictive judgments based on that perception. This intuitive alignment between visual understanding and temporal inference has inspired researchers to transform time series data into image-like representations, such as spectrograms, recurrence plots, or frequency-enhanced textures (Ni et al. 2025), and directly leverage pre-trained vision models for downstream time series analysis tasks. Representative examples include TimesNet (Wu et al. 2023a), which recasts sequences into two-dimensional formats for CNN processing; VisionTS (Chen et al. 2024), fine-tuning Masked Autoencoders (MAE) on visualized time series (He et al. 2022); and TimeVLM (Zhong et al. 2025), leveraging vision-language models for multimodal forecasting. These approaches repurpose vision models’ capability to detect edges, gradients, and frequency textures for identifying temporal structures and dynamics.
@@
However, repurposing large vision models (LVMs) for TSF introduces substantial redundancy and misalignment. These models are originally architected for image data rich in semantic content – a modality that fundamentally diverges from the purely numerical signals of time series. To investigate this misalignment and understand which visual features benefit time series forecasting, we conduct t-SNE visualization of features extracted by MAE from ImageNet (Deng et al. 2009) and four time-series benchmarks (ECL, Weather, Electricity, ETT), as shown in Figure 1. The results reveal that while some visual features exhibit overlap and similarity with time series data, others show significant distributional differences between the two modalities. Fig.1 (a) and (b) lie at the interface of temporal and visual features, exhibiting similar texture patterns, while Fig.1 (c) and (d) appear in distant image clusters, demonstrating richer semantic content. Further analysis uncovers a distinct binary differentiation pattern: time-series features align closely with texture-rich images such as metal meshes and stage curtains, yet diverge markedly from semantically complex scenes like waterbirds and parachuting activities. This pattern suggests that vision models capture low-level textural features relevant to time series, while their high-level semantic representations not only add unnecessary complexity but also impair forecasting by introducing overfitting to irrelevant visual features. In limited-data scenarios, this redundancy becomes particularly detrimental as models latch onto spurious visual patterns instead of essential temporal dynamics.
@@ 2 Related Work
Transforming time series data into visual representations has emerged as a popular paradigm in forecasting research. However, existing approaches have yet to systematically investigate which visual features genuinely benefit forecasting performance and which introduce harmful noise. Early efforts converted sequences into images and trained convolutional neural networks (CNNs) from scratch to perform forecasting tasks (Wang et al. 2015; Sezer et al. 2018; Li et al. 2020; Sood et al. 2021; Semenoglou et al. 2023).
As this line of research has progressed, structural-level and model-level innovations have both played key roles. On the structural side, TimesNet decomposes time series into multi-periodic components and rearranges them into 2D tensors (Wu et al. 2023b), while TimeMixer and its successor TimeMixer++ apply mixing operations across time and feature dimensions to capture patterns at multiple scales (Wang et al. 2024a, b). Meanwhile, model-level adaptations harness the power of vision backbones pre-trained on natural images: BEiT was first repurposed for forecasting (Zhou et al. 2023), VisionTS demonstrated the effectiveness of Masked Autoencoders (MAE) on grayscale time-series images (Chen et al. 2024), TimeVLM introduced multimodal forecasting with vision-language models (Zhong et al. 2025), and LDM4TS leveraged multi-view diffusion models for enhanced prediction (Ruan et al. 2025).
@@ 3 Methodology
Teacher-Student Model Module. Implements an asymmetric design where a large pre-trained vision model (teacher) first learns to identify beneficial cross-modal patterns, then transfers this knowledge to a lightweight encoder (student). The teacher generates high-quality supervisory signals during training and guides the student through knowledge distillation. After training, only the efficient student model is retained for deployment.
Knowledge Distillation Module. Transfers predictive knowledge through pyramid-style feature alignment and selective attention distillation, ensuring the student replicates only the most forecasting-relevant behaviors. Adaptive loss weighting dynamically balances effective knowledge transfer with optimal prediction accuracy.
@@ 3.1 Cross-Modal Representation Module
Temporal Feature Extraction. Given an input sequence , we employ a patch-based transformer architecture. The sequence is segmented into overlapping patches of length , each projected into a -dimensional embedding space with positional encodings. These embeddings are processed through transformer layers:
where . The final representation captures multi-scale hierarchical temporal patterns essential for accurate forecasting.
Visual Augmentation for Time Series. Our analysis reveals that time series align with texture-rich visual patterns rather than semantic content. This motivates a transformation pipeline that emphasizes low-level visual features:
@@
where , FFT extracts essential frequency components and PE adds trigonometric encodings with period to precisely capture temporal cycles.
Multi-Scale Transformation. The augmented features undergo hierarchical convolutions. First, 1D to capture fine-grained local dependencies, then 2D layers to create spatial patterns. This multi-scale approach effectively captures diverse patterns from local variations to global trends.
@@
This visual representation reveals temporal patterns invisible in 1D sequences while providing a format compatible with pre-trained vision models, specifically targeting the minimal vision capabilities useful for time series analysis.
@@ 3.2 Teacher-Student Model Design
Visual Feature Extraction. The teacher employs large pre-trained vision backbones to extract visual features:
where and is the fusion dimension for aligning different modalities. Despite containing predominantly redundant parameters for time series tasks, these large models enable the teacher to explore which visual patterns genuinely benefit forecasting.
@@
The fused representation combines attention output with temporal features:
@@
where represents predictions for future time steps across variables. To ensure high-quality supervisory signals, we optimize:
Beyond predictions, the teacher provides attention matrices capturing temporal dependencies, fused representations , and soft prediction targets for distillation.
| Methods | Ours | Only Teacher | Only Student | Time-VLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | FEDformer | Autoformer | |||||||||||||
| Methods | Ours | Only Teacher | Only Student | Time-VLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | FEDformer | Autoformer | |||||||||||||
@@ 4.1 Experimental Settings
Datasets & Evaluation Metrics. We evaluate OccamVTS on eight benchmark datasets: ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, Traffic (Zhou et al. 2021; Lai et al. 2018), and M4 (Makridakis et al. 2018). Performance is measured using MAE and MSE for the first seven datasets, while M4 uses SMAPE, MASE, and OWA following competition protocols (Oreshkin et al. 2019). Dataset details and metric specifications are in Appendix A and B.
Baselines. We extensively compare OccamVTS with state-of-the-art time series models and ablation variants (teacher-only and student-only configurations), including recent vision-augmented methods like TimeVLM (Zhong et al. 2025), LDM4TS (Ruan et al. 2025), TimeMixer++ (Wang et al. 2024b), TimeMixer (Wang et al. 2024a), and TimesNet (Wu et al. 2023a); transformer-based architectures like iTransformer (Liu et al. 2023), PatchTST (Nie et al. 2022), FEDformer (Zhou et al. 2022), ETSformer (Woo et al. 2022), Non-Stationary Transformer (Liu et al. 2022b), Autoformer (Wu et al. 2021), and Informer (Zhou et al. 2021); and highly competitive linear models like DLinear (Zeng et al. 2023) and LightTS (Campos et al. 2023).
Implementation Details. We use a unified benchmarking pipeline (Wu et al. 2023a) with a knowledge distillation framework. The teacher model can be selected from six pre-trained vision architectures (MAE variants, CLIP, EfficientNet-B3, ResNet-101), while the student model can be chosen from three lightweight architectures (EfficientNet-B0, MobileNet, Tiny-ViT). Models are trained using Adam optimizer with learning rate on NVIDIA RTX A6000 GPU (48GB). See Appendix C for details.
| Methods | Ours | Only Teacher | Only Student | Times-VLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | Autoformer | ||||||||||||
@@
4.2 Long-term Forecasting
4.3 Few-shot Forecasting
To evaluate the data efficiency and generalization capability of our proposed method, we conducted comprehensive few-shot learning experiments using only 10% of the training data. As shown in Table 2, our method achieved the best performance on five out of seven datasets (ETTh1, ETTh2, ETTm1, ETTm2, and Weather) in terms of both MSE and MAE metrics. On the ETTh1 dataset, we obtain 2.1% improvement over TimeVLM, while on ETTm2, our approach achieves significant 8.2% MSE reduction compared to the second-best performer. The ablation experiments in the few-shot setting further validate the importance of the knowledge distillation mechanism, where the complete method consistently outperforms the variant without knowledge distillation by 3.0-4.6% across all datasets. This indicates that the knowledge distillation component is particularly valuable when training data is limited, as it enables more effective cross-modal knowledge transfer from pre-trained vision models to lightweight temporal architectures.
4.4 Zero-shot Forecasting
We conduct zero-shot transfer experiments across ETT datasets without any fine-tuning to evaluate cross-domain generalization capabilities. As shown in Table 3, OccamVTS achieves the best performance in 5 MSE and 7 MAE metrics out of 8 scenarios, demonstrating strong cross-domain transferability. For challenging transfer tasks like ETTh2->ETTh1 and ETTm2->ETTm1, our approach achieves 13.5% and 6.7% MSE improvements over TimesVLM respectively, significantly outperforming other baseline methods. The ablation study confirms that knowledge distillation consistently improves zero-shot performance across all transfer scenarios, with particularly notable improvements of 5.4% on ETTh2->ETTh1 and 10.7% on ETTm2->ETTm1 when comparing KD versus non-KD variants. While most baseline methods show significant performance degradation (20-40%) in cross-dataset transfers, our method maintains consistent performance with only ±8% variation across different transfer pairs, suggesting robust generalization capabilities through effective distillation of universal temporal patterns from vision models.

### v2

@@ Abstract
Time series forecasting is fundamental to diverse applications, with recent approaches leveraging large vision models (LVMs) to capture temporal patterns through visual representations. We reveal that while vision models enhance forecasting performance, 99% of their parameters are unnecessary for time series tasks. Through cross-modal analysis, we find that time series align with low-level textural features but not high-level semantics, which can impair forecasting accuracy. We propose OccamVTS, a knowledge distillation framework that extracts only the essential 1% of predictive information from LVMs into lightweight networks. Using pre-trained LVMs as privileged teachers, OccamVTS employs pyramid-style feature alignment combined with correlation and feature distillation to transfer beneficial patterns while filtering out semantic noise. Counterintuitively, this aggressive parameter reduction improves accuracy by eliminating overfitting to irrelevant visual features while preserving essential temporal patterns. Extensive experiments across multiple benchmark datasets demonstrate that OccamVTS consistently achieves state-of-the-art performance with only 1% of the original parameters, particularly excelling in few-shot and zero-shot scenarios.
Code — https://github.com/sisuolv/OccamVTS
@@ 1 Introduction
Time series forecasting (TSF) is a fundamental task in machine learning, underpinning a wide range of critical applications, including energy demand prediction, financial market analysis, meteorological modeling, and traffic flow optimization (idrees2019prediction; kiyasseh2021clocs; xu2021anomaly; bi2023accurate). The objective of TSF is to anticipate future values based on historical observations of one or more temporally evolving variables. Despite its broad utility, TSF presents significant challenges due to inherent characteristics such as non-stationarity, long-range dependencies, stochastic noise, and the simultaneous presence of localized patterns and global trends across multiple temporal scales.
@@
Deep learning has emerged as the dominant paradigm for TSF, with state-of-the-art models consistently achieving strong performance across diverse benchmarks. These methods operate directly on numerical sequences or their frequency-domain counterparts, enabling rich representations of complex temporal dynamics. Convolutional neural networks (CNNs) and Transformers, for example, are particularly adept at capturing both short-term fluctuations and long-term dependencies (ismail2019deep); advanced models such as Autoformer (wu2021autoformer) and FEDformer (zhou2022fedformer) explicitly decompose sequences into trend and seasonal components (bai2018empirical; hatami2018classification; li2019logtrans; liu2022pyraformer; wen2023transformers). Distinct from other modalities that rely on abstract semantics, these architectures effectively model time series as structured numerical signals, ultimately grounding predictions in precise and interpretable temporal behaviors.
@@
In parallel, vision models have recently gained significant traction as an alternative and intriguing approach to TSF, driven by the observation that humans can often see meaningful patterns in time series plots. For instance, a seasoned bond trader may visually recognize a shift in trend or momentum in a price curve and form predictive judgments based on that perception. This intuitive alignment between visual understanding and temporal inference has inspired researchers to transform time series data into image-like representations, such as spectrograms, recurrence plots, or frequency-enhanced textures (ni2025harnessing), and directly leverage pre-trained vision models for downstream time series analysis tasks. Representative examples include TimesNet (wu2022timesnet), which recasts sequences into two-dimensional formats for CNN processing; VisionTS (chen2024visionts), fine-tuning Masked Autoencoders (MAE) on visualized time series (he2022masked); and TimeVLM (zhong2025time), leveraging vision-language models for multimodal forecasting. These approaches repurpose vision models’ capability to detect edges, gradients, and frequency textures for identifying temporal structures and dynamics.
@@
However, repurposing large vision models (LVMs) for TSF introduces substantial redundancy and misalignment. These models are originally architected for image data rich in semantic content – a modality that fundamentally diverges from the purely numerical signals of time series. To investigate this misalignment and understand which visual features benefit time series forecasting, we conduct t-SNE visualization of features extracted by MAE from ImageNet (deng2009imagenet) and four time-series benchmarks (ECL, Weather, Electricity, ETT), as shown in Figure 1. The results reveal that while some visual features exhibit overlap and similarity with time series data, others show significant distributional differences between the two modalities. Fig.1 (a) and (b) lie at the interface of temporal and visual features, exhibiting similar texture patterns, while Fig.1 (c) and (d) appear in distant image clusters, demonstrating richer semantic content. Further analysis uncovers a distinct binary differentiation pattern: time-series features align closely with texture-rich images such as metal meshes and stage curtains, yet diverge markedly from semantically complex scenes like waterbirds and parachuting activities. This pattern suggests that vision models capture low-level textural features relevant to time series, while their high-level semantic representations not only add unnecessary complexity but also impair forecasting by introducing overfitting to irrelevant visual features. In limited-data scenarios, this redundancy becomes particularly detrimental as models latch onto spurious visual patterns instead of essential temporal dynamics.
@@ 2 Related Work
Transforming time series data into visual representations has emerged as a popular paradigm in forecasting research. However, existing approaches have yet to systematically investigate which visual features genuinely benefit forecasting performance and which introduce harmful noise. Early efforts converted sequences into images and trained convolutional neural networks (CNNs) from scratch to perform forecasting tasks (wang2015encoding; sezer2018algorithmic; li2020forecasting; sood2021visual; semenoglou2023image).
As this line of research has progressed, structural-level and model-level innovations have both played key roles. On the structural side, TimesNet decomposes time series into multi-periodic components and rearranges them into 2D tensors (wu2023timesnet), while TimeMixer and its successor TimeMixer++ apply mixing operations across time and feature dimensions to capture patterns at multiple scales (wang2024timemixer; wang2024timemixer++). Meanwhile, model-level adaptations harness the power of vision backbones pre-trained on natural images: BEiT was first repurposed for forecasting (zhou2023one), VisionTS demonstrated the effectiveness of Masked Autoencoders (MAE) on grayscale time-series images (chen2024visionts), TimeVLM introduced multimodal forecasting with vision-language models (zhong2025time), and LDM4TS leveraged multi-view diffusion models for enhanced prediction (ruan2025vision).
@@ 3 Methodology
Teacher-Student Model Module. Implements an asymmetric design where a frozen pretrained LVM serves as teacher with only its fusion and prediction heads trained to produce privileged supervisory signals. The lightweight student encoder is jointly trained through forecasting and distillation losses to learn cross-modal patterns from the teacher. After training, only the efficient student model is retained for deployment.
Knowledge Distillation Module. Transfers predictive knowledge through pyramid-style feature alignment and selective attention distillation, ensuring the student replicates only forecasting-relevant behaviors. Correlation distillation aligns temporal attention while feature distillation employs a composite objective to match fused representations. Adaptive loss weighting dynamically balances knowledge transfer with prediction accuracy.
@@ 3.1 Cross-Modal Representation Module
Temporal Feature Extraction. Given an input time series , where denotes a batch of multivariate sequences, is batch size, sequence length, and the number of variables. We employ a patch-based transformer architecture. The sequence is segmented into overlapping patches of length with stride , each projected into a -dimensional embedding space with positional encodings. These embeddings are processed through transformer layers:
where , is the patch embedding sequence, and is the number of encoder blocks. We then pool the patch tokens to obtain a sequence-level embedding , where is the number of temporal patches. This pooled embedding serves as the temporal query in cross-modal fusion.
Visual Augmentation for Time Series. Our analysis reveals that time series data align with texture-rich visual patterns rather than high-level semantic content. This motivates a transformation pipeline that emphasizes low-level visual features to exploit pre-trained vision models:
@@
where stacks the raw window, frequency, and periodic channels for each length- window, and we instantiate the two operators as
Here denotes the normalized length- window used for spectral analysis, is the time index within a window, the frequency index, the transform length, and the dataset-specific periodicity hyperparameter (Appendix A). In practice, we use the normalized magnitude of .
Multi-Scale Transformation. The augmented features undergo hierarchical convolutions. First, lightweight depthwise–pointwise 1D blocks to capture short-range local dependencies, then shallow 2D layers on the (channeltime) grid to create spatial patterns. This yields intermediate tensors before resizing, enabling local variations and global trends to interact as visual textures that pre-trained vision encoders are biased to detect.
@@
where denote image height/width (in pixels) after resizing and should not be confused with the forecasting horizon . Concretely, for a target pixel we use bilinear interpolation
where are the four nearest neighbors and are distance-based weights; min-max normalization then scales intensities to the vision backbone’s range
This interpolation-normalization pipeline aligns pseudo-images with pre-trained vision encoders, revealing temporal patterns invisible in 1D sequences while leveraging minimal vision capabilities necessary for time series analysis.
@@ 3.2 Teacher-Student Model Design
Visual Feature Extraction. The teacher employs frozen large pre-trained vision backbones to extract visual features:
where removes spatial dimensions, projects the backbone output to the fusion dimension, and is the fusion dimension for aligning different modalities. We deliberately employ global pooling to prevent noisy spatial alignment between synthetic images and natural-image priors, extracting only aggregate visual features relevant to temporal patterns. Despite containing predominantly redundant parameters for time series tasks, these large models enable the teacher to explore which visual patterns genuinely benefit forecasting.
@@
Here and are learned projections, and is the key/query width. The fused representation combines attention output with temporal features:
@@
where , is the forecasting horizon, and is the number of predicted variables. To ensure high-quality supervisory signals, we optimize:
where denotes ground‑truth targets. Beyond predictions, the teacher provides attention matrices capturing temporal dependencies, fused representations , and soft prediction targets for distillation.
@@ 3.3 Knowledge Distillation Module
3.4 Contributions at a glance
Overall, our methodological innovations are threefold: we design a texture-oriented cross-modal representation that converts time series into pseudo-images through frequency and periodicity enhancement and a lightweight 1D2D multi-scale transformation, followed by spatial reorganization and global pooling to robustly couple with the low-level inductive biases of pre-trained vision backbones; we propose an asymmetric teacher–student architecture that freezes the vision teacher and performs cross-attention fusion where temporal queries attend to aggregated visual keys and values to form privileged representations; and we introduce a scalable selective knowledge transfer scheme that combines pyramid-style feature alignment with two complementary objectives, namely correlation distillation on temporal attention and distillation at the representation level, so that the student preserves forecasting relevant textures while suppressing semantic redundancy and noise.
@@ 4 Experiments
| Methods | Ours | Only Teacher | Only Student | TimeVLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | FEDformer | Autoformer | |||||||||||||
| Methods | Ours | Only Teacher | Only Student | TimeVLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | FEDformer | Autoformer | |||||||||||||
@@ 4.1 Experimental Settings
Datasets & Evaluation Metrics. We evaluate OccamVTS on eight benchmark datasets: ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, Traffic (zhou2021informer; lai2018modeling), and M4 (makridakis2018m4). Performance is measured using MAE and MSE for the first seven datasets, while M4 uses SMAPE, MASE, and OWA following competition protocols (oreshkin2019n). Dataset details and metric specifications are in Appendix A and B.
Baselines. We extensively compare OccamVTS with state-of-the-art time series models and ablation variants (teacher-only and student-only configurations), including recent vision-augmented methods like TimeVLM (zhong2025time), LDM4TS (ruan2025vision), TimeMixer++ (wang2024timemixer++), TimeMixer (wang2024timemixer), and TimesNet (wu2022timesnet); transformer-based architectures like iTransformer (liu2023itransformer), PatchTST (nie2022time), FEDformer (zhou2022fedformer), ETSformer (woo2022etsformer), Non-Stationary Transformer (liu2022non), Autoformer (wu2021autoformer), and Informer (zhou2021informer); and highly competitive linear models like DLinear (zeng2023transformers) and LightTS (campos2023lightts).
Implementation Details. We use a unified benchmarking pipeline (wu2022timesnet) with a knowledge distillation framework. From the teacher pool (MAE variants, CLIP, EfficientNet-B3, ResNet-101) and the student pool (EfficientNet-B0, MobileNet-V3, Tiny-ViT), we instantiate one teacher and one student per experiment; unless otherwise noted, we adopt MAE-Large as the teacher and Tiny-ViT as the student, which corresponds to 1% of the teacher’s total parameters. Models are trained using Adam optimizer with learning rate on NVIDIA RTX A6000 GPU (48GB). See Appendix C for details.
4.2 Long-term Forecasting
4.3 Few-shot Forecasting
To evaluate the data efficiency and generalization capability of our proposed method, we conducted comprehensive few-shot learning experiments using only 10% of the training data. As shown in Table 2, our method achieved the best performance on five out of seven datasets (ETTh1, ETTh2, ETTm1, ETTm2, and Weather) in terms of both MSE and MAE metrics. On the ETTh1 dataset, we obtain 2.1% improvement over TimeVLM, while on ETTm2, our approach achieves significant 4% MSE reduction compared to the second-best performer. The ablation experiments in the few-shot setting further validate the importance of the knowledge distillation mechanism, where the complete method consistently outperforms the variant without knowledge distillation by 3.0-4.6% across all datasets. This indicates that the knowledge distillation component is particularly valuable when training data is limited, as it enables more effective cross-modal knowledge transfer from pre-trained vision models to lightweight temporal architectures.
4.4 Zero-shot Forecasting
We conduct zero-shot transfer experiments across ETT datasets without any fine-tuning to evaluate cross-domain generalization capabilities. As shown in Table 3, OccamVTS achieves the best performance in 5 MSE and 7 MAE metrics out of 8 scenarios, demonstrating strong cross-domain transferability. For challenging transfer tasks like ETTh2->ETTh1 and ETTm2->ETTm1, our approach achieves 13.5% and 6.7% MSE improvements over TimeVLM respectively, significantly outperforming other baseline methods. The ablation study confirms that knowledge distillation consistently improves zero-shot performance across all transfer scenarios, with particularly notable improvements of 5.4% on ETTh2->ETTh1 and 10.7% on ETTm2->ETTm1 when comparing KD versus non-KD variants. While most baseline methods show significant performance degradation (20-40%) in cross-dataset transfers, our method maintains consistent performance with only ±8% variation across different transfer pairs, suggesting robust generalization capabilities through effective distillation of universal temporal patterns from vision models.
| Methods | Ours | Only Teacher | Only Student | TimeVLM | TimeMixer++ | TimeMixer | LDM4TS | TimesNet | iTransformer | DLinear | PatchTST | Autoformer | ||||||||||||